Company DELTA uses perpetual inventory method and its weighted-average method. The company commenced operations on January 1, 2018, and purchased 400 units of inventory, with each unit costing $ 7.5. On January 10, 2018, the company purchased an additional 600 units, each unit costing ISK. 9.0. On January 11, 2018, the Company sold 550 units of the inventory.

What would be the position on the company's inventory account immediately after the sale?

A. 3.080

B. 3.780

C. 3.713

D. 4.738
Answer #1

Weighted average cost per unit

= Total cost/Total units

= [(400*7.5)+(600*9)]/ (400+600) = 8400/1000 = 8.40 per unit

Units in ending inventory = (400+600) - 550 = 450 units

Cost of inventory = 450 * 8.40

= 3780

Option B